\begin{thebibliography}{99}
\small
\addtolength{\itemsep}{-1.5 em}
\setlength{\itemsep}{-5pt}
\bibitem{Alic}M. Ali\'{c}, P.S. Bullen, J. Pe\v{c}ari\'{c}, V. Volenec, On the geometric-arithmetic mean inequality for matrices, Math. Commun. 2 (1997) 125-128.
\bibitem{Ando1}T. Ando, C.K. Li, R. Mathias, Geometric means, Linear Algebra Appl. 385 (2004) 305-334.
\bibitem{Ando2}T. Ando, Concavity of certain maps on positive definite matrices and applications to Hadamard products, Linear Algebra Appl. 26 (1979) 203-241.
\bibitem{Ando3}T. Ando, Topics on operator inequalities, Hokkaido Univ. Lecture Note, 1978.
\bibitem{Bhatia}R. Bhatia, Positive definite matrices, Princeton (NJ): Princeton University Press; 2007.
\bibitem{Bhatia2}R. Bhatia, F. Kittaneh, Notes on matrix arithmetic-geometric mean inequalities, Linear Algebra Appl. 308 (2000) 203-211.
\bibitem{Bhatia3}R. Bhatia, R.L. Karandikar, Monotonicity of the matrix geometric mean, Math. Ann. 353 (2012) 1453-1467.
\bibitem{Fujii} J.I. Fujii, M. Fujii, M. Nakamura, J. Pe\v{c}ari\'{c}, Y. Seo, A reverse inequality for the weighted geometric mean due to Lawson-Lim, Linear Algebra Appl.  427 (2007) 272-284.
\bibitem{Fujii1}J.I. Fujii, Y. Seo, On the Ando-Li-Mathias mean and the Karcher mean of positive definite matrices, Linear Multilinear Algebra, 63 (2015) 636-649.
\bibitem{Fujii2}M. Fujii, S. Izumino, R. Nakamoto, Y. Seo, Operator inequalities related to Cauchy-Schwarz and H\"{o}lder-McCarthy inequalities, Nihonkai Math. J. 8 (1997) 117-122.
\bibitem{Furuta1}T. Furuta, Operator inequalities associated with H\"{o}lder-McCarthy and Kantorovich inequalities, J. Inequal. Appl. 2 (1998) 137-148.
\bibitem{Kubo}F. Kubo, T. Ando, Means of positive linear operators, Math. Ann. 246 (1980) 205-224.
\bibitem{Lawson2}J. Lawson, Y. Lim, A general framework for extending means to higher orders. Colloq. Math. 113 (2008) 191-221.
\bibitem{Lawson3}J. Lawson, Y. Lim, Karcher means and Karcher equations of positive definite operators, Trans. Amer. Math. Soc., Series b, 1 (2014) 1-22.
\bibitem{Lim} Y. Lim, M. P\'{a}lfia, Matrix power means and the Karcher mean, Journal of Functional Analysis, 262 (2012) 1498-1514.
\bibitem{Lin1}M. Lin, Squaring a reverse AM-GM inequality, Studia Math. 215 (2013) 187-194.
\bibitem{Lin2}M. Lin, On an operator Kantorovich inequality for positive linear maps, J. Math. Anal. Appl. 402 (2013) 127-132.
\bibitem{Micic} J. Mi\'{c}i\'{c}, J. Pe\v{c}ari\'{c}, Yuki Seo, Complementary inequalities to inequalities of Jensen and Ando based on the Mond-Pe\v{c}ari\'{c} method, Linear Algebra Appl. 318 (2000) 87-107.
\bibitem{Moakher} M. Moakher, A differential geometric approch to the geometric mean of symmetric positive-definite matrices, SIAM J. Matrix Anal. Appl. 26 (2005) 735-747.
\bibitem{Pusz} W. Pusz, S.L. Woronowicz, Functional calculus for sesquilinear forms and the purification map, Rep. Math. Phys. 8 (1975) 159-170.
\bibitem{Seo}Y. Seo, Generalized P\'{o}lya-Szeg\"{o} type inequalities for some non-commutative geometric means, Linear Algebra Appl. 438 (2013) 1711-1726.
\bibitem{Sra} S. Sra, R. Hosseini, Conic geometric optimisation on the manifold of positive definite matrices, arXiv:1312.1039v3.
\bibitem{Tominaga} M. Tominaga, Specht's ratio in the Young inequality, Sci. Math. Japon. 55 (2002) 583-588.
\bibitem{Wang} S.G. Wang, M.X. Wu, Z.Z. Jia, Matrix inequalities, Beijing: Science Press, 2006.
\bibitem{Yamazaki}T. Yamazaki, An extension of Kantorovich inequality to $n$-operators via the geometric mean by Ando-Li-Mathias, Linear Algebra Appl. 416 (2006) 688-695.


%\bibitem{Alic}M. Ali\'{c}, P.S. Bullen, J. Pe\v{c}ari\'{c}, V. Volenec. On the geometric-arithmetic mean inequality for matrices. Math. Commun. 1997;2:125-128.
%\bibitem{Ando1}T. Ando, C.K. Li, R. Mathias, Geometric means. Linear Algebra Appl. 2004; 385: 305-334.
%\bibitem{Ando2}T. Ando, Concavity of certain maps on positive definite matrices and applications to Hadamard products, Linear Algebra Appl. \textbf{26} (1979), 203-241.
%\bibitem{Ando3}T. Ando, Topics on operator inequalities, Hokkaido Univ. Lecture Note, 1978.
%\bibitem{Bhatia}R. Bhatia, Positive definite matrices, Princeton (NJ): Princeton University Press; 2007.
%%\bibitem{Bhatia1}R. Bhatia and C. Davis, More operator versions of the Schwarz inequality, Comm. Math. Phys. \textbf{215} (2000), 239-244.
%\bibitem{Bhatia2}R. Bhatia and F. Kittaneh, Notes on matrix arithmetic-geometric mean inequalities, Linear Algebra Appl. \textbf{308} (2000), 203-211.
%\bibitem{Bhatia3}R. Bhatia, R.L. Karandikar, Monotonicity of the matrix geometric mean, Math. Ann. (2012) 353:1453-1467.
%%\bibitem{Furuta2}T. Furuta, J. Mi\'{c}i\'{c} Hot, J. Pe\v{c}ari\'{c} and Y. Seo, Mond-Pe\v{c}ari\'{c} Method in Operator Inequalities, Element, Zagreb, 2005.
%%\bibitem{Fu1}X. Fu and C. He, Some operator inequalities for positive linear maps, Linear Multilinear Algebra, \textbf{63}(3) (2015), 571-577.
%%\bibitem{Fur2} S. Furuichi, Refined Young inequalities with Specht's ratio, J. Egypt. Math. Soc.,  20 (2012), 46--49.
%%\bibitem{Fu2}X. Fu, Some generalizations of operator inequalities, Preprint.
%\bibitem{Fujii} J.I. Fujii, M. Fujii, M. Nakamura, J. Pe\v{c}ari\'{c}, Y. Seo, A reverse inequality for the weighted geometric mean due to Lawson-Lim, Linear Algebra Appl.  427 (2007) 272-284.
%\bibitem{Fujii1}J.I. Fujii, Y. Seo, On the Ando-Li-Mathias mean and the Karcher mean of positive definite matrices, Linear Multilinear Algebra, 63(3)(2015), 636-649.
%\bibitem{Fujii2}M. Fujii, S. Izumino, R. Nakamoto, Y. Seo, Operator inequalities related to Cauchy-Schwarz and H\"{o}lder-McCarthy inequalities, Nihonkai Math. J. 8(1997) 117-122.
%\bibitem{Furuta1}T. Furuta, Operator inequalities associated with H\"{o}lder-McCarthy and Kantorovich inequalities, J. Inequal. Appl. 2 (1998) 137-148.
%%\bibitem{Gumus}I. H. Gumus, A note on a conjecture about Wielandt's inequality, Linear Multilinear Algebra, (2014), doi:10.1080/03081087.2014.982496.
%%\bibitem{Hirzallah1} O. Hirzallah, F. Kittaneh, M. Krni\'{c}, N. Lovri\v{c}evi\'{c} and J. Pe\v{c}ari\'{c}, Refinements and reverses of means inequalities for Hilbert space operators, Banach J. Math. Anal. \textbf{7} (2013), no. 2, 15--29.
%%\bibitem{Hoa} D. T. Hoa, D. T. H. Binh and H. M. Toan, On some inequalities with means. Preprint.
%%\bibitem{Hardy} G.H. Hardy, J.E. Little-wood and G. P\'{o}lya, Inequalities, Cambridge at the university press, 1934.
%%\bibitem{Horn}R.A. Horn and C.R. Johnson, Matrix analysis, 2nd ed, Cambridge University Press, New York, 2013.
%%\bibitem{Kantorovich}L.V. Kantorovich, Functional analysis and applied mathematics (in Russian), Uspehi Matem. Nauk (N.S.). \textbf{3} (1948), 89-185.
%\bibitem{Kubo}F. Kubo and T. Ando, Means of positive linear operators, Math. Ann. \textbf{246} (1980), 205-224.
%%\bibitem{Liao}W. Liao, J. Wu and J. Zhao, New versions of reverse Young and heinz mean inequalities with the Kantorovich constant, Taiwanese J. Math. (2014). doi: 10.11650/tjm.18.2014.4548.
%\bibitem{Lawson2}J. D. Lawson, Y. Lim. A general framework for extending means to higher orders. Colloq. Math. 2008; 113: 191-221.
%\bibitem{Lawson3}J. D. Lawson and Y. Lim, Karcher means and Karcher equations of positive definite operators. Trans. Amer. Math. Soc., Series b, 1(2014), 1-22.
%\bibitem{Lim} Y. Lim, M. P\'{a}lfia, Matrix power means and the Karcher mean, Journal of Functional Analysis, 262 (2012) 1498-1514.
%\bibitem{Lin1}M. Lin, Squaring a reverse AM-GM inequality, Studia Math. \textbf{215} (2013), 187-194.
%\bibitem{Lin2}M. Lin, On an operator Kantorovich inequality for positive linear maps, J. Math. Anal. Appl. \textbf{402} (2013), 127-132.
%%\bibitem{Marshall} A.W. Marshall and I. Olkin, Matrix versions of the Cauchy and Kantorovich inequalities, Aequationes Math. \textbf{40} (1990), 89-93.
%\bibitem{Micic} J. Mi\'{c}i\'{c}, J. Pe\v{c}ari\'{c}, Yuki Seo, Complementary inequalities to inequalities of Jensen and Ando based on the Mond-Pe\v{c}ari\'{c} method, Linear Algebra Appl.318 (2000) 87-107.
%\bibitem{Moakher} M. Moakher. A differential geometric approch to the geometric mean of symmetric positive-definite matrices. SIAM J. Matrix Anal. Appl. 2005;26:735-747.
%\bibitem{Pusz} W. Pusz, S.L. Woronowicz, Functional calculus for sesquilinear forms and the purification map, Rep. Math. Phys. 8 (1975) 159-170.
%%\bibitem{Roberts} A.W. Roberts and D.E. Varberg, Convex Functions, Academic Press, New York, 1973.
%\bibitem{Seo}Y. Seo, Generalized P\'{o}lya-Szeg\"{o} type inequalities for some non-commutative geometric means, Linear Algebra Appl. 438 (2013) 1711-1726.
%\bibitem{Sra} S. Sra, R. Hosseini, Conic geometric optimisation on the manifold of positive definite matrices, arXiv:1312.1039v3.
%\bibitem{Tominaga} M. Tominaga, Specht's ratio in the Young inequality, Sci. Math. Japon. \textbf{55} (2002), 583-588.
%%\bibitem{Yamazaki} T. Yamazaki, M. Yanagida, Characterizations of chaotic order associated with Kantorovich inequality, Sci. Math. 2 (1) (1999) 37-50.
%\bibitem{Wang} S.G. Wang, M.X. Wu, Z.Z. Jia, Matrix inequalities, Beijing: Science Press, 2006.
%\bibitem{Yamazaki}T. Yamazaki, An extension of Kantorovich inequality to $n$-operators via the geometric mean by Ando-Li-Mathias, Linear Algebra Appl. 416 (2006) 688-695.
%%\bibitem{Zhang} P. Zhang, More operator inequalities for positive linear maps, Banach J. Math. Anal. \textbf{9} (1) (2015), 166-172.
%%\bibitem{Zuo} H. Zuo, G. Shi and M. Fujii, Refined Young inequality with Kantorovich constant, J. Math. Inequal., 5 (2011), 551-556.
%%\bibitem{Lawson1}Jimmie D. Lawson and Yongdo Lim, The geometric mean, matrices, metrics, and more, Amer. Math. Monthly 108 (2001), no. 9, 797-812.


\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{color}
\usepackage{graphicx}
\usepackage{enumerate}
\usepackage{amsfonts,amsmath,latexsym,amssymb}
\usepackage{amsthm}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{corollary}{Corollary}[section]
\newtheorem{lemma}{Lemma}[section]
\newtheorem{proposition}{Proposition}[section]
\newtheorem{remark}{Remark}[section]
\newtheorem{definition}{Definition}[section]

\numberwithin{equation}{section}

\begin{document}
\title{On the power means and Lawson-Lim means for positive invertible operators}
\author{Wenshi Liao$^a$\footnote{Corresponding author. E-mail: liaowenshi@gmail.com, jlwu678@163.com, 527628144@qq.com.}, Junliang Wu$^a$ and Haisong Cao$^a$\\
$^a${\small College of Mathematics and Statistics, Chongqing University, Chongqing 401331, P.R. China}}

\date{}
\maketitle
{\bf Abstract.} This note aims to
present some reverse inequalities about the power means and Karcher mean via the Kantorovich constant and some of these have been generalized to higher power. Also, we generalize the reverse weighted arithmetic-geometric mean inequality of $n$ positive invertible operators due to Lawson and Lim. In addition, we make comparisons between the Karcher mean and Lawson-Lim geometric mean for higher power.
\vspace{3mm}

{\bf Keywords : } Power means; Karcher mean; Lawson-Lim geometric mean; Ando-Li-Mahthias geometric mean; Kantorovich constant; Reverse inequalities
\vspace{3mm}


{\bf AMS Subject Classification :} 47A30; 47A63; 47A64
\vspace{3mm}

\section{Introduction}
Let $B(\mathcal{H})$ be the $C^*$-algebra of all bounded linear operators on a complex separable Hilbert space $\mathcal{H}$. $B(\mathcal{H})^+$ stands for the set of positive elements in $B(\mathcal{H})$.  A linear map $\Phi$: $B(\mathcal{H})\rightarrow B(\mathcal{K})$ is said to be positive if $\Phi(A)\ge 0$ whenever $A\ge0$.
A positive linear map is said to be normalized (unital) if it maps $I_n$, the identity operator, to $I_m$.
Note that a positive linear map $\Phi$ is monotone in the sense that $A\le B$ implies $\Phi(A)\le \Phi(B)$.
 $\mathbb{P}$ stands for the convex cone of positive invertible operators. $\Delta_n$ denotes the simplex of positive probability vectors in $\mathbb{R}^n$ convexly spanned by the unit coordinate vectors. $\|\cdot\|$ and $|||\cdot|||$ denote the operator norm and the unitarily invariant norm, respectively. $\mathrm{tr}$ is the trace functional.

In \cite{Lim}, Lim and P\'{a}lfia have proposed power means of positive definite matrices
 and their notion and most of their results readily extended to the setting of positive invertible operators on a complex Hilbert space (See \cite{Lawson2}):
 \begin{definition}{\rm(\textbf{Power means})}
Let $\mathbb{A}=(A_1,A_2,\cdots,A_n)\in \mathbb{P}^n$ and $\omega\in \Delta_n$.
For $t\in[-1,1]$,
the power means $P_t(\omega;\mathbb{A})$ is defined as the unique
positive definite solution of the following non-linear equations:
\[
X=\sum\limits^n_{i=1}\omega_i(X\sharp_tA_i), ~for~ t\in(0,1],~~~~~
\]
\[
~~~~~~X=\sum\limits^n_{i=1}\omega_i(X^{-1}\sharp_{-t}A_i^{-1})^{-1}, ~for~ t\in[-1,0),
\]
where $A\sharp_t B :=A^{\frac{1}{2}}(A^{-\frac{1}{2}}BA^{-\frac{1}{2}})^{t} A^{\frac{1}{2}}$
is the $t$-weighted geometric mean of $A$ and $B$. $P_t(\omega;\mathbb{A})$ is called
the $\omega$-weighted power mean of order $t$ of $A_1,A_2,\cdots,A_n$.  To simplify the notation,
we write $P_t(\mathbb{A})=P_t(\frac{1}{n},\frac{1}{n},\cdots,\frac{1}{n};\mathbb{A})$.
\end{definition}

Since the power means $P_t(\omega;\mathbb{A})$ is increasing for $t\in [-1,1]\setminus \{0\}$,
 it interpolates between the weighted arithmetic mean and harmonic mean: $P_1(\omega;\mathbb{A})=\sum_{i=1}^nw_iA_i$
and $P_{-1}(\omega;\mathbb{A})=(\sum_{i=1}^nw_iA_i^{-1})^{-1}$.

Let $t\in(0,1]$ and $f:\mathbb{P}\rightarrow \mathbb{P}$ defined by $f(X)=\sum^n_{i=1}\omega_i(X\sharp_tA_i)$.
Then by the L\"{o}ewner-Heinz inequality, $f$ is monotone: $X\le Y$ implies $f(X)\le f(Y)$.
By Theorem 3.1 and Remark 3.4 of \cite{Lawson3}, $f$ is a strict contraction for the Thompson metric and
has a unique fixed point by the Banach fixed point theorem:
\[\lim_{k\rightarrow\infty}f^k(X)=P_t(\omega;\mathbb{A}), ~X\in\mathbb{P}.\]

Since the pioneering papers of Pusz and Woronowicz \cite{Pusz}, Ando \cite{Ando3}, and Kubo and Ando \cite{Kubo},
an extensive theory of two-variable geometric mean has sprung up for positive operators:
For two positive operators $A$ and $B$, the operator geometric mean is defined by
$A\sharp B :=A^{\frac{1}{2}}(A^{-\frac{1}{2}}BA^{-\frac{1}{2}})^{\frac{1}{2}} A^{\frac{1}{2}}$
for $A>0$.  Once one realizes that the matrix geometric mean $A\sharp B$ is the metric midpoint of $A$ and $B$
for the trace metric on the set of positive definite matrices of some fixed dimension (see, e.g., \cite{Bhatia, Lawson2}).
The operator geometric mean has many characterizations. For example,
\[A\sharp B=\max\left\{X\mid X=X^*,
\left[
  \begin{array}{cc}
    A  & X  \\
    X & B \\
  \end{array}
\right]\ge0
\right\}
\]
and $A\sharp B$ is the unique positive solution of the Riccati equation $XA^{-1}X = B$. Moreover, it is monotone,
jointly concave and congruence invariant and $A\sharp B=B\sharp A$.
But the $n$-variable case for $n> 2$ was a long standing problem and many authors studied the geometric mean of $n$-variable.

In 2004, Ando et al.\cite{Ando1} succeeded in the formulate of the geometric mean for $n$ positive definite matrices,
and they showed that it satisfies ten important properties:
\begin{definition}
{\rm(\textbf{Ando-Li-Mahthias geometric mean}\cite{Ando1})}
Let $A_i$ $(i=1,2,\cdots,n)$ be positive definite matrices.
Then the geometric mean $G_{ALM}(A_1,A_2,\cdots,A_n)$ is defined by induction as follows:

$~\mathrm{(i)}$ $G_{ALM}(A_1,A_2)=A_1\#A_2$.

$\mathrm{(ii)}$ Assume that the geometric mean of any $n-1$-tuple of operators is defined. Let
\[
G_{ALM}((A_j)_{j\neq i})=G_{ALM}(A_1,\cdots,A_{i-1},A_{i+1},\cdots,A_n),
\]
\indent \indent\and let sequences $\{A^{(r)}_i\}^\infty_{r=0}$ be $A^{(0)}_i=A_i$ and $A^{(r)}_i =G_{ALM}((A^{(r-1)}_j)_{j\neq i})$. If there exists
\indent \indent $\lim_{r\rightarrow \infty}A^{(r)}_i$, and it does not depend on $i$, then the geometric mean of $n$-matrices is
\indent \indent defined as
\[\lim\limits_{r\rightarrow \infty}A^{(r)}_i=G_{ALM}(A_1,A_2,\cdots,A_n).\]
\end{definition}
In [15], Yamazaki pointed out that the definition of the geometric mean by Ando, Li and
Mathias can be extended to Hilbert space operators.
Lawson and Lim \cite{Lawson2} established a definition of the weighted version of
the Ando-Li-Mahthias geometric mean for $n$ positive operators, we call it \textbf{Lawson-Lim geometric mean}. Following \cite{Lawson2},
we recall the definition of higher order weighted geometric mean $G[n,t]$ with $t\in(0,1)$ for $n$
positive operators $A_1,A_2,\cdots,A_n$. Let
$G[2,t](A_1,A_2)=A_1\sharp_t A_2 =A_1^{\frac{1}{2}}(A_1^{-\frac{1}{2}}A_2A_1^{-\frac{1}{2}})^{t} A_1^{\frac{1}{2}}$
(the unique geodesic curve containing $A$ and $B$ and its unique metric midpoint
 $A\sharp B=A\sharp_{\frac{1}{2}}B$ is the geometric mean of $A$ and $B$).
For $n\ge 3$, $G[n,t]$ is defined inductively as follows: Put $A_i^{(0)}=A_i$ for all $i=1,2,\cdots,n$ and
\[
A_i^{(r)}=G[n-1,t]\left(\left(A_j^{(r-1)}\right)_{j\neq i}\right)=G[n-1,t]\left(A_1^{(r-1)},\cdots,A_{i-1}^{(r-1)},A_{i+1}^{(r-1)},\cdots,A_n^{(r-1)}\right)
\]
inductively for $r$. Then the sequences $\{A_i^{(r)}\}$ have the same limit for all $i=1,2,\cdots,n$ in the Thompson metric.
So $G[n,t](A_1,A_2,\cdots,A_n)=\lim_{r\rightarrow\infty}A_i^{(r)}$.
In particular, $G[n,\frac{1}{2}]$ for $t=\frac{1}{2}$ is the Ando-Li-Mathias geometric mean.

Similarly, the weighted arithmetic mean is defined as follows: Put $\widetilde{A_i^{(0)}}=A_i$ for all $i=1,2,\cdots,n$ and
\[
\widetilde{A_i^{(r)}}=A[n-1,t]\left(\left(\widetilde{A_j^{(r-1)}}\right)_{j\neq i}\right)=A[n-1,t]\left(\widetilde{A_1^{(r-1)}},\cdots,\widetilde{A_{i-1}^{(r-1)}},\widetilde{A_{i+1}^{(r-1)}},\cdots,\widetilde{A_n^{(r-1)}}\right)
\]
inductively for $r$. Then the sequences $\{\widetilde{A_i^{(r)}}\}$ have the same limit for all $i=1,2,\cdots,n$.
If we put $A[n,t](A_1,A_2,\cdots,A_n)=\lim_{r\rightarrow\infty}\widetilde{A_i^{(r)}}$, then it is expressed by
\[A[n,t](A_1,A_2,\cdots,A_n)=t[n]_1A_1+t[n]_2A_2+\cdots+t[n]_nA_n,\]
where $t[n]_i\ge0$ for all $i=1,2,\cdots,n$ with $\sum_{i=1}^n t[n]_i=1$. Also, the weighted harmonic mean
$H[n,t](A_1,A_2,\cdots,A_n)$ is defined as
\[H[n,t](A_1,A_2,\cdots,A_n)=\left(t[n]_1A_1^{-1}+t[n]_2A_2^{-1}+\cdots+t[n]_nA_n^{-1}\right)^{-1}.\]
Note that the coefficient $\{t[n]_i\}$ depends on $n$ and $t$ only, see \cite{Fujii, Seo} for more details.

Moreover, the weighted arithmetic-geometric-harmonic mean inequality holds:
\begin{equation}
\label{agh}
H[n,t](A_1,A_2,\cdots,A_n)\le G[n,t](A_1,A_2,\cdots,A_n) \le A[n,t](A_1,A_2,\cdots,A_n).
\end{equation}
Since then, another approach to generalizing the geometric mean to $n$-variables, depending on Riemannian trace metric, was the Karcher mean, which was studied by many researchers, see \cite{Lawson3, Lim} and the reference therein.
Let $\mathbb{A}=A_1, A_2,\cdots,A_n\in \mathbb{P}^n$ and $\omega=(w_1,w_2,\cdots, w_n)\in\Delta_n$. By computing appropriate derivatives as in \cite{Bhatia, Moakher},
 the $\omega$-weighted \textbf{Karcher mean} of $\mathbb{A}$, denoted by $G_K(\omega;\mathbb{A})$,
coincides with the unique positive definite solution of the Karcher equation
\begin{equation}
\label{Kar}
\sum_{i=1}^nw_i\log(X^{\frac{1}{2}}A_i^{-1}X^{-\frac{1}{2}})=0.
\end{equation}
In the two operators case, $A_1, A_2\in \mathbb{P}$, the Karcher mean coincides with the weighted geometric mean $A_1\sharp_t A_2 =A_1^{\frac{1}{2}}(A_1^{-\frac{1}{2}}A_2A_1^{-\frac{1}{2}})^{t} A_1^{\frac{1}{2}}$.
From \eqref{Kar}, the Karcher mean satisfies the self-duality
$G_K(\omega;\mathbb{A})=G_K(\omega;\mathbb{A}^{-1})^{-1},$
where $\mathbb{A}^{-1}=(A^{-1}_1,A^{-1}_2,\cdots,A^{-1}_n)$.
Lim and P\'{a}lfia \cite{Lim} also established that the Karcher mean is the limit of power means as $t\rightarrow0$ in the finite-dimensional setting:
\begin{equation}
\label{Lim}
\lim_{t\rightarrow0} P_t(\omega;\mathbb{A})=G_K(\omega;\mathbb{A}),
\end{equation}
then Lawson and Lim \cite{Lawson3} showed \eqref{Lim} is valid in the infinite-dimensional setting.

\section{Preliminary}

Tominaga \cite{Tominaga} and Ali\'{c} et al.\cite{Alic} showed the following non-commutative arithmetic-geometric mean inequality:
For positive invertible operators $A_1$ and $A_2$ such that $m\le A_1, A_2\le M$ for some scalars $0<m\le M$ and $h=\frac{M}{m}$, one has
\begin{equation}\label{tom}
(1-v)A_1+vA_2\le S(h)A_1\sharp_vA_2,
\end{equation}
where $S(h)=\frac{(h-1)h^{\frac{1}{h-1}}}{e\log h}(h\neq1)$ is the Specht ratio and $S(1)=1$.
By \eqref{tom}, we can easily obtain the following one for any positive unital linear map $\Phi$:
\begin{equation}\label{tom1}
\Phi(A_1)\sharp_v\Phi(A_2)\le (1-v)\Phi(A_1)+v\Phi(A_2)=\Phi((1-v)A_1+vA_2)\le S(h)\Phi(A_1\sharp_v A_2).
\end{equation}

Fujii et al.\cite{Fujii} showed the following reverses of the weighted
arithmetic-geometric mean inequality of $n$ positive invertible operators due to Lawson and Lim:
\begin{theorem}\label{th21}
For any integer $n\ge 2$, let $A_1,A_2,\cdots,A_n$ be positive invertible operators
such that $m\le A_i\le M$ for all $i=1,2,\cdots,n$ and some scalars $0<m\le M$.
Then for each $t\in(0,1)$
\begin{equation}\label{th211}
~~A[n,t](A_1,\cdots, A_n)\le \frac{(m+M)^{2}}{4mM}G[n,t](A_1,\cdots, A_n)
\end{equation}
and
\begin{equation}\label{th212}
A[n,t](A_1,\cdots, A_n)\le S^2(h)G[n,t](A_1,\cdots, A_n).~~~
\end{equation}
\end{theorem}
Fujii \cite{Fujii1} and Lin \cite{Lin1} state a relation between the Specht ratio and the Kantorovich constant severally:
For $0<m\le M$ and $h=\frac{M}{m}$,
\begin{equation}\label{SK}
S(h)\le \frac{(m+M)^{2}}{4mM}\le S^2(h).
\end{equation}

It is obvious that \eqref{th211} is tighter than \eqref{th212} via the inequality \eqref{SK}.

Seo \cite{Seo} presented the following inequalities for the $n$-variables in terms of the higher order
weighted geometric mean due to Lawson and Lim, which are an extension and a converse of Ando's inequality
 $\Phi(A\sharp_t B)\le \Phi(A)\sharp_t \Phi(B)$
(See \cite[Theorem 4.1.5]{Bhatia} and \cite{Ando2}) for a positive unital linear map $\Phi$ and $t\in[0,1]$, respectively:
\begin{theorem}\label{th22}
Let $\Phi$ be a positive unital linear map on $B(\mathcal{H})$ and
let $A_1, A_2, \cdots, A_n \in \mathbb{P}^n$ for any positive integer $n\ge 2$ on a Hilbert space $\mathcal{H}$ such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$.
Then for each $t\in(0,1)$,
\[
\Phi(G[n,t](A_1,\cdots, A_n))\le G[n,t](\Phi(A_1),\cdots, \Phi(A_n))\le \frac{(m+M)^{2}}{4mM}\Phi(G[n,t](A_1,\cdots, A_n)).
\]
\end{theorem}
On the other hand, Bhatia \cite{Bhatia3} derive an inequality of the $\omega$-weighted Karcher mean of $n$ positive definite matrices for any positive unital linear map $\Phi$ that
\begin{equation}\label{Bhatia}
\Phi(G_K(\omega;A_1,\cdots, A_n))\le G_K(\omega;\Phi(A_1),\cdots, \Phi(A_n)).
\end{equation}
Also, Lim \cite{Lim} obtained the above inequality via the power means. We will be engaged to show the reverse inequality of \eqref{Bhatia}.

Although the Ando-Li-Mahthias geometric mean does not coincide with the Karcher mean in general, Fujii and Seo \cite{Fujii1} made a comparison between the Ando-Li-Mahthias geometric mean and the Karcher mean and also obtained an inequality for the Ando-Li-Mahthias geometric mean:
\begin{theorem} \cite{Fujii1}
For any integer $n\ge 2$, let $A_1,A_2,\cdots,A_n$ be positive definite matrices in $\mathbb{P}$
such that $m\le A_i\le M$ for all $i=1,2,\cdots,n$ and some scalars $0<m\le M$.
Then
\begin{equation}\label{th231}
\frac{4mM}{(m+M)^{2}}G_K(A_1,\cdots, A_n)\le G_{ALM}(A_1,\cdots, A_n)\le\frac{(m+M)^{2}}{4mM}G_K(A_1,\cdots, A_n)~~~
\end{equation}
and
\begin{equation}\label{th232}
G_{ALM}(A^p_1,\cdots,A^p_n)\le \frac{(m+M)^{2p}}{4^pm^pM^p}G_{ALM}(A_1,\cdots, A_n)^p~~for~all~0<p<1.
\end{equation}
\end{theorem}
The case of $p>1$ about \eqref{th232} will be one of motivations for our study.

\section{The power means and Karcher mean}
It is well known that the Power means and Karcher mean are relevant via \eqref{Lim}. In this section, we will obtain some generalized properties and inequalities with respect to the Power means and Karcher mean.

Firstly, we list some important properties of the power means, some of which appeared in \cite {Lawson3, Lim}:
\begin{proposition}\label{pro}
Let $\mathbb{A}=(A_1,\cdots, A_n)\in \mathbb{P}^n$, $\mathbb{A}^{-1}=(A^{-1}_1,A^{-1}_2,\cdots,A^{-1}_n) \in \mathbb{P}^n$, a weight vector $\omega=(w_1,\cdots,w_n)\in \Delta_n$
and $t\in[-1,1]\setminus \{0\}$.
Then the power means satisfies the following properties:

 $(\mathrm{P1})$ (Duality) $P_{-t}(\omega;\mathbb{A}^{-1})^{-1}=P_{t}(\omega;\mathbb{A});$

$(\mathrm{P2})$ (Homogeneity) $P_t(\omega;a\mathbb{A})=a P_t(\omega;\mathbb{A});$

 $(\mathrm{P3})$ (Continuous) $\lim_{t\rightarrow 0}P_t(\omega;\mathbb{A})=G_K (\omega;\mathbb{A});$

 $(\mathrm{P4})$ ( APH weighted mean inequalities)  $(\sum_{i=1}^nw_iA_i^{-1})^{-1}\le P_t(\omega;\mathbb{A})\le \sum_{i=1}^nw_iA_i;$

 $(\mathrm{P5})$ If $t\in(0,1]$, then $\Phi(P_t(\omega;\mathbb{A}))\le P_t(\omega;\Phi(\mathbb{A}))$ for any positive unital linear map $\Phi$,
where
\indent\ \indent\ $\Phi(\mathbb{A})=(\Phi(A_1),\cdots,\Phi(A_n))$.
If $t\in [-1,0)$, then $P_t(\omega;\Phi(\mathbb{A}))\le\Phi(P_t(\omega;\mathbb{A}))$ for any
\indent\ \indent\ strictly positive unital linear map $\Phi;$

$(\mathrm{P6})$ For $t\in(0,1]$,
\[
\mathrm{tr}\left(P_t(\omega;\mathbb{A})\right)\le \left(\sum\limits_{i=1}^nw_i(\mathrm{tr}A_i)^t\right)^{\frac{1}{t}} ~and~
\mathrm{tr}\left(P_{-t}(\omega;\mathbb{A})\right)\ge n\left(\sum\limits_{i=1}^nw_i(\mathrm{tr}A^{-1}_i)^t\right)^{-\frac{1}{t}}.
\]
\end{proposition}
\begin{proof}
We provide a proof of $(\mathrm{P6})$. The other properties are known in \cite{Lawson3, Lim}.

By the definition of the power means, let $X=P_t(\omega;\mathbb{A})$ for $t\in(0,1]$. Then
\[\mathrm{tr}X=\sum_{i=1}^nw_i\mathrm{tr}(X\sharp_tA_i)\le \sum_{i=1}^nw_i(\mathrm{tr}X)^{1-t}(\mathrm{tr}A_i)^t=(\mathrm{tr}X)^{1-t}\sum_{i=1}^nw_i(\mathrm{tr}A_i)^t,\]
where the inequality follows from the Corollary 9 of \cite{Sra}.
Thus, $\mathrm{tr}X\le \left(\sum_{i=1}^nw_i\mathrm{tr}(A_i)^t\right)^{\frac{1}{t}}$.

Since
$n=\mathrm{tr}I\le \mathrm{tr}A \cdot\mathrm{tr}A^{-1}$ for an $n$-dimensional operator $A> 0$ (See \cite[Theorem 6.2.2]{Wang}),
\[\mathrm{tr}(P_{-t}(\omega;\mathbb{A}))=\mathrm{tr}(P_{t}(\omega;\mathbb{A}^{-1})^{-1})\ge n \left(\mathrm{tr}P_{t}(\omega;\mathbb{A}^{-1})\right)^{-1}\ge n\left(\sum\limits_{i=1}^nw_i(\mathrm{tr}A^{-1}_i)^t\right)^{-\frac{1}{t}}.\qedhere\]
\end{proof}

\begin{remark}
Let $\Phi(A)=\frac{\mathrm{tr}A}{n}$ for an $n$-dimensional operator $A> 0$ in $(\mathrm{P5})$ when $t\in(0,1]$ or by the first inequality in Proposition \ref{pro}{\rm (P6)}.  As $t\rightarrow 0$, by \eqref{Lim}, then we have
\[
\mathrm{tr}(G_K(\omega;\mathbb{A}))\le \prod_{i=1}^n(\mathrm{tr}A_i)^{w_i}.
\]
\end{remark}
\begin{remark}
Let $\Phi(A)=\frac{\mathrm{tr}A}{n}$ in the Choi's inequality $\Phi(A)^{-1}\le \Phi(A^{-1})$ for an $n$-dimensional operator $A> 0$ (See \cite[Theorem 2.3.6]{Bhatia}). Then we have $\mathrm{tr}A^{-1}\ge n^2\left(\mathrm{tr}A\right)^{-1}.$
 So we can obtain a stronger result than the second inequality in {\rm (P6)} as follows:
 \[
 \mathrm{tr}(P_{-t}(\omega;\mathbb{A}))\ge n^2\left(\sum\limits_{i=1}^nw_i(\mathrm{tr}A^{-1}_i)^t\right)^{-\frac{1}{t}}.
\]
\end{remark}

The next Proposition generalize properties (P6).

\begin{proposition}\label{th32}
Let $A_1, A_2, \cdots, A_n\in \mathbb{P}^n$ be $n$-dimensional operators such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then for $t\in(0,1]$,
\begin{equation}\label{trace1}
\mathrm{tr}(P_t(\omega;\mathbb{A}))\ge \left(\frac{4mM}{(m+M)^2} \sum\limits_{i=1}^nw_i(\mathrm{tr}A_i)^t\right)^{\frac{1}{t}}~~~~~
\end{equation}
and
\[
~~~~~~~~~~~~\mathrm{tr}(P_{-t}(\omega;\mathbb{A}))\le n^2\cdot \left(\frac{(m+M)^2}{4mM}\right)^{1+\frac{1}{t}}\left(\sum\limits_{i=1}^nw_i(\mathrm{tr}A^{-1}_i)^t\right)^{-\frac{1}{t}}.
\]
\end{proposition}

\begin{proof}
By Theorem \ref{th22} with $n=2$, the following inequality
\[
\Phi(A)\sharp_t\Phi(B)\le \frac{(m+M)^2}{4mM}\Phi(A\sharp_t B)
\]
holds. Let $\Phi(A)=\frac{\mathrm{tr}A}{n}$ for $A>0$ in the above inequality and $X=P_t(\omega;\mathbb{A})$ for $t\in(0,1]$ in the definition of power means. Then
\begin{align*}
\mathrm{tr}X&=\sum_{i=1}^nw_i\mathrm{tr}(X\sharp_tA_i)\\&\ge \frac{4mM}{(m+M)^2}\sum_{i=1}^nw_i(\mathrm{tr}X)^{1-t}(\mathrm{tr}A_i)^t\\
&=\frac{4mM}{(m+M)^2}(\mathrm{tr}X)^{1-t}\sum_{i=1}^nw_i(\mathrm{tr}A_i)^t,
\end{align*}
therefore,
\[
(\mathrm{tr}X)^t\ge \frac{4mM}{(m+M)^2}\sum_{i=1}^nw_i(\mathrm{tr}A_i)^t.
\]
Taking $\Phi(A)=\frac{\mathrm{tr}A}{n}$ in the Kantorovich inequality $\Phi(A^{-1})\le\frac{(m+M)^2}{4mM}\Phi(A)^{-1}$ for any $A > 0$(See \cite[Proposition 2.7.8]{Bhatia}), we have
\[\mathrm{tr}A^{-1}\le \frac{n^2(m+M)^2}{4mM}(\mathrm{tr}A)^{-1}~ \mathrm{for}~ A> 0.\]
Hence,
\begin{align*}
\mathrm{tr}(P_{-t}(\omega;\mathbb{A}))&=\mathrm{tr}(P_{t}(\omega;\mathbb{A}^{-1})^{-1})\\
&\le \frac{n^2(m+M)^2}{4mM} (\mathrm{tr}(P_{t}(\omega;\mathbb{A}^{-1})))^{-1}\\
&\le n^2\left(\frac{(m+M)^2}{4mM}\right)^{1+\frac{1}{t}}\left(\sum\limits_{i=1}^nw_i(\mathrm{tr}A^{-1}_i)^t\right)^{-\frac{1}{t}}.\qedhere
\end{align*}
\end{proof}
\begin{remark}
By \eqref{tom1} and a similar way as in the proof of Proposition \ref{th32}, we derive
\[
\mathrm{tr}(P_t(\omega;\mathbb{A}))\ge \left(S(h)^{-1} \sum\limits_{i=1}^nw_i(\mathrm{tr}A_i)^t\right)^{\frac{1}{t}}.
\]
From \eqref{SK}, we know that the above inequality is sharper than \eqref{trace1}.
\end{remark}

\begin{theorem}\label{th31}
Let $\Phi$ be a positive unital linear map on $B(\mathcal{H})$ and
let $A_1, A_2, \cdots, A_n \in \mathbb{P}^n$ for any positive integer $n\ge 2$ on a Hilbert space $\mathcal{H}$ such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then for $t\in(0,1]$,
\begin{equation}
\label{th311}
\sum\limits_{i=1}^nw_i\Phi(A_i)\le \frac{(m+M)^2}{4mM}\Phi(P_t(\omega;\mathbb{A}))~~~~~~
\end{equation}
and
\[
~~~~~~~P_t(\omega;\Phi(\mathbb{A}))\le\frac{(m+M)^2}{4mM}\left(\sum\limits_{i=1}^nw_i\Phi(A_i)^{-1}\right)^{-1}.
\]
\end{theorem}
\begin{proof}
By the Kantorovich inequality (See \cite[Proposition 2.7.8]{Bhatia}) and the APH weighted mean inequalities $(\mathrm{P4})$, we can easily obtained that
\[
\sum\limits_{i=1}^nw_i\Phi(A_i)=\Phi\left(\sum\limits_{i=1}^nw_iA_i\right)\le \frac{(m+M)^2}{4mM} \Phi\left(\left(\sum\limits_{i=1}^nw_iA_i^{-1}\right)^{-1}\right)\le\frac{(m+M)^2}{4mM}\Phi(P_t(\omega;\mathbb{A}))
\]
and
\[P_{t}(\omega;\Phi(\mathbb{A}))\le \sum\limits_{i=1}^nw_i\Phi(A_i)\le\frac{(m+M)^2}{4mM} \left(\sum\limits_{i=1}^nw_i\Phi(A_i)^{-1}\right)^{-1}\label{ph}.\qedhere\]
\end{proof}

Note that Theorem \ref{th31} generalize the APH weighted mean inequalities (P4) and by \eqref{Lim}, the Karcher mean satisfies the AKH weighted mean inequalities (See \cite{Lim})
\begin{equation}
\label{akh}
\left(\sum_{i=1}^nw_iA_i^{-1}\right)^{-1}\le G_K(\omega;\mathbb{A})\le \sum_{i=1}^nw_iA_i.
\end{equation}

By \eqref{Lim} and Theorem \ref{th31}, we generalize the AKH weighted mean inequalities \eqref{akh} as follows:
\begin{corollary}
 Under the same conditions as in Theorem \ref{th31}, then
\begin{equation}
\label{agk}
\sum\limits_{i=1}^nw_i\Phi(A_i)\le \frac{(m+M)^2}{4mM} \Phi(G_K(\omega;\mathbb{A}))~~~~~
\end{equation}
and
\[
~~~~~G_K(\omega;\Phi(\mathbb{A}))\le\frac{(m+M)^2}{4mM} \left(\sum\limits_{i=1}^nw_i\Phi(A_i)^{-1}\right)^{-1}\label{ph}.
\]
\end{corollary}

\begin{theorem}\label{th32}
Let $\Phi$ be a positive unital linear map on $B(\mathcal{H})$ and
let $A_1, A_2, \cdots, A_n \in \mathbb{P}^n$ for any positive integer $n\ge 2$ on a Hilbert space $\mathcal{H}$ such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then
\begin{equation}\label{th321}
P_t(\omega;\Phi(\mathbb{A}))\le \frac{(m+M)^2}{4mM}\Phi(P_t(\omega;\mathbb{A})) ~for ~t\in (0,1]~~
\end{equation}
and
\[
~\Phi(P_t(\omega;\mathbb{A}))\le \frac{(m+M)^2}{4mM}P_t(\omega;\Phi(\mathbb{A}))~for ~t\in [-1,0).
\]
\end{theorem}

\begin{proof}
For $t\in (0,1]$, by the weighted APH mean inequalities and \eqref{th311}, it follows that
\[P_t(\omega;\Phi(\mathbb{A}))\le \sum_{i=1}^nw_i\Phi(A_i)\le \frac{(m+M)^2}{4mM} \Phi(P_t(\omega;\mathbb{A})).\]
Let $t \in [-1,0)$ and let $\Phi$ be a strictly positive unital linear map. By the Kantorovich inequality (the reverse of Choi's inequality)
(See \cite[Proposition 2.7.8]{Bhatia}), $\Phi(A^{-1})\le\frac{(m+M)^2}{4mM}\Phi(A)^{-1}$ for all $A > 0$ and \eqref{th321}, we have
\[
k P_{-t}(\omega;\Phi(\mathbb{A})^{-1})=P_{-t}(\omega;k\Phi(\mathbb{A})^{-1})\le P_{-t}(\omega;\Phi(\mathbb{A}^{-1}))\le k\Phi(P_{-t}(\omega;\mathbb{A}^{-1})),
\]
where $k=\frac{(m+M)^2}{4mM}$ and $\Phi(\mathbb{A})^{-1}=(\Phi(A_1)^{-1},\cdots, \Phi(A_n)^{-1})$. Therefore,
\[P_{-t}(\omega;\Phi(\mathbb{A})^{-1})\le \Phi(P_{-t}(\omega;\mathbb{A}^{-1})).\]
This implies that
\[
\Phi(P_{-t}(\omega;\mathbb{A}^{-1})^{-1})\le k\Phi(P_{-t}(\omega;\mathbb{A}^{-1}))^{-1}
\le k P_{-t}(\omega;\Phi(\mathbb{A})^{-1})^{-1}.
\]
By $(\mathrm{P1})$, we obtain the desired inequality
\[
\Phi(P_t(\omega; \mathbb{A}))\le\frac{(m+M)^2}{4mM}P_t(\omega; \Phi(\mathbb{A})).\qedhere
\]
\end{proof}

By \eqref{Lim} and Theorem \ref{th32}, we show a reverse inequality of \eqref{Bhatia} for $n$ positive invertible operators as follows:
\begin{corollary}
 Under the same conditions as in Theorem \ref{th32}, then
\[
G_K(\omega;\Phi(\mathbb{A}))\le \frac{(m+M)^2}{4mM}\Phi(G_K(\omega;\mathbb{A})).
\]
\end{corollary}

Next, we are devoted to obtain several higher power inequalities which are related to
inequalities \eqref{th311} and \eqref{agk}. In order to do that, we need two important lemmas.

\begin{lemma}\cite[Lemma 2.1]{Bhatia2}
\label{lema1}
Let $A, B\ge 0$. Then the following inequality holds:
\begin{equation}
\label{lem12}
\|AB\|\le\frac{1}{4}\|A+B\|^2.
\end{equation}
\end{lemma}

\begin{lemma}\cite[p. 28]{Bhatia}
\label{lema2}
Let $A, B\ge 0$. Then for $1\le r<+\infty$,
\begin{equation}
\label{lem2}
\|A^r+B^r\|\le\|(A+B)^r\|.
\end{equation}
\end{lemma}

It is well known that $\|A\|\le 1$ is equivalent to $A^*A\le I$. This fact plays an important role in the proof of theorems.
\begin{theorem}\label{th33}
Let $\Phi$ be a positive unital linear map on $B(\mathcal{H})$ and
let $A_1, A_2, \cdots, A_n \in \mathbb{P}^n$ for any positive integer $n\ge 2$ on a Hilbert space $\mathcal{H}$ such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then for $t\in(0,1]$ and $p\ge 2$,
\begin{equation}
\label{ap1}
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(m+M)^{2p}}{16m^pM^p} \Phi(P_t(\omega;\mathbb{A}))^p
\end{equation}
and
\begin{equation}
\label{ap2}
~\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(m+M)^{2p}}{16m^pM^p} P_t(\omega;\Phi(\mathbb{A}))^p.
\end{equation}
\end{theorem}

\begin{proof}
It is known that \eqref{ap1} is equivalent to
\[
\|\Phi(\sum\limits_{i=1}^nw_iA_i)^\frac{p}{2}\Phi(P_t(\omega;\mathbb{A}))^{-\frac{p}{2}}\|\le \frac{(m+M)^{p}}{4m^\frac{p}{2}M^\frac{p}{2}} .
\]
If $p\ge2$, then 
\begin{align*}
\|M^\frac{p}{2}m^\frac{p}{2}&\Phi(\sum\limits_{i=1}^nw_iA_i)^\frac{p}{2}\Phi(P_t(\omega;\mathbb{A}))^{-\frac{p}{2}}\|\\
&\le \frac{1}{4}\|\Phi(\sum\limits_{i=1}^nw_iA_i)^\frac{p}{2}+M^\frac{p}{2}m^\frac{p}{2}\Phi(P_t(\omega;\mathbb{A}))^{-\frac{p}{2}}\|^2~~(\mathrm{by}~ \eqref{lem12})\\
&\le \frac{1}{4}\|\Phi(\sum\limits_{i=1}^nw_iA_i)+Mm\Phi(P_t(\omega;\mathbb{A}))^{-1}\|^p~~(\mathrm{by}~ \eqref{lem2})\\
&\le \frac{1}{4}\|\Phi(\sum\limits_{i=1}^nw_iA_i)+Mm\Phi(P_t(\omega;\mathbb{A})^{-1})\|^p~~(\text{by the Choi's inequality})\\
&\le \frac{1}{4}\|\Phi(\sum\limits_{i=1}^nw_iA_i)+Mm\Phi(\sum\limits_{i=1}^nw_iA_i^{-1})\|^p~~(\mathrm{by}~ (\rm{P}4))\\
&\le \frac{(M+m)^p}{4}.
\end{align*}
The last inequality above holds as follows:  The condition $0<m\le A_i\le M$ implies that
\[A_i+MmA_i^{-1}\le M+m.\]
Therefore, the following inequality
\begin{equation}
\label{app}
\Phi(\sum\limits_{i=1}^nw_iA_i)+Mm\Phi(\sum\limits_{i=1}^nw_iA_i^{-1})\le M+m
\end{equation}
holds for any positive unital linear map $\Phi$ and $\omega=(w_1,\cdots,w_n)\in \Delta_n$. Thus, \eqref{ap1} holds.

The inequality \eqref{ap2} is equivalent to
\[
\|\Phi(\sum\limits_{i=1}^nw_iA_i)^\frac{p}{2}P_t(\omega;\Phi(\mathbb{A}))^{-\frac{p}{2}}\|\le \frac{(m+M)^{p}}{4m^\frac{p}{2}M^\frac{p}{2}} .
\]
To prove \eqref{ap2}, by computing, we have
\begin{align*}
\|M^\frac{p}{2}m^\frac{p}{2}&\Phi(\sum\limits_{i=1}^nw_iA_i)^\frac{p}{2}P_t(\omega;\Phi(\mathbb{A}))^{-\frac{p}{2}}\|\\
&\le \frac{1}{4}\|\Phi(\sum\limits_{i=1}^nw_iA_i)^\frac{p}{2}+M^\frac{p}{2}m^\frac{p}{2}P_t(\omega;\Phi(\mathbb{A}))^{-\frac{p}{2}}\|^2~~(\mathrm{by}~ \eqref{lem12})\\
&\le \frac{1}{4}\|\Phi(\sum\limits_{i=1}^nw_iA_i)+MmP_t(\omega;\Phi(\mathbb{A}))^{-1}\|^p~~(\mathrm{by}~ \eqref{lem2})\\
&\le \frac{1}{4}\|\Phi(\sum\limits_{i=1}^nw_iA_i)+Mm\sum\limits_{i=1}^nw_i\Phi(A_i)^{-1}\|^p~~(\mathrm{by}~ (\rm{P}4))\\
&\le \frac{1}{4}\|\Phi(\sum\limits_{i=1}^nw_iA_i)+Mm\Phi(\sum\limits_{i=1}^nw_iA_i^{-1})\|^p~~(\text{by the Choi'sinequality})\\
&\le \frac{(M+m)^p}{4}.~~(\mathrm{by}~ \eqref{app})
\end{align*}
Thus, \eqref{ap2} holds.
\end{proof}

By \eqref{Lim}, we obtain the following results about the Karcher mean and the arithmetic mean for any positive unital linear map.
\begin{corollary}\label{cor33}
 Under the same conditions as in Theorem 3.1, then
\[
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(m+M)^{2p}}{16m^pM^p} \Phi(G_K(\omega;\mathbb{A}))^p
\]
and
\[
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(m+M)^{2p}}{16m^pM^p} G_K(\omega;\Phi(\mathbb{A}))^p.
\]
\end{corollary}

By Theorem \ref{th33}, Corollary \ref{cor33} and the L\"{o}ewner-Heinz inequality, we have
\begin{remark}
For $p\in(0,2]$, we have
\begin{equation}
\label{rem331}
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(m+M)^{2p}}{4^pm^pM^p} \Phi(P_t(\omega;\mathbb{A}))^p,
\end{equation}
\begin{equation}
\label{rem332}
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(m+M)^{2p}}{4^pm^pM^p}P_t(\omega;\Phi(\mathbb{A}))^p,
\end{equation}
\begin{equation}
\label{rem351}
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(m+M)^{2p}}{4^pm^pM^p} \Phi(G_K(\omega;\mathbb{A}))^p,
\end{equation}
and
\[
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(m+M)^{2p}}{4^pm^pM^p} G_K(\omega;\Phi(\mathbb{A}))^p.
\]
\end{remark}

\begin{theorem}\label{th34}
Let $\Phi$ be a positive unital linear map on $B(\mathcal{H})$ and
let $A_1, A_2, \cdots, A_n \in \mathbb{P}^n$ for any positive integer $n\ge 2$ on a Hilbert space $\mathcal{H}$ such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then for $t\in(0,1]$, $1<\alpha\le 2$ and $p\ge2\alpha$,
\begin{equation}
\label{th341}
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(k^{\frac{\alpha}{2}}(M^\alpha+m^\alpha))^{\frac{2p}{\alpha}}}{16M^pm^p} \Phi(P_t(\omega;\mathbb{A}))^p
\end{equation}
and
\begin{equation}
\label{th342}
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(k^{\frac{\alpha}{2}}(M^\alpha+m^\alpha))^{\frac{2p}{\alpha}}}{16M^pm^p} P_t(\omega;\Phi(\mathbb{A}))^p.
\end{equation}
\end{theorem}
\begin{proof}
It follows from the inequality \eqref{rem331} that
\begin{equation}
\label{th343}
\Phi(P_t(\omega;\mathbb{A}))^{-\alpha}\le k^\alpha\Phi(\sum\limits_{i=1}^nw_iA_i)^{-\alpha} ~\text{for}~ 1<\alpha\le 2.
\end{equation}
It is known that \eqref{th341} is equivalent to
\[
\|[\Phi(\sum\limits_{i=1}^nw_iA_i)^\frac{p}{2}\Phi(P_t(\omega;\mathbb{A}))^{-\frac{p}{2}}\|\le \frac{(k^{\frac{\alpha}{2}}(M^\alpha+m^\alpha))^{\frac{p}{\alpha}}}{4M^{\frac{p}{2}}m^{\frac{p}{2}}}.
\]
If $p\ge2\alpha$, then 
\begin{align*}
\|M^\frac{p}{2}m^\frac{p}{2}&\Phi(\sum\limits_{i=1}^nw_iA_i)^\frac{p}{2}\Phi(P_t(\omega;\mathbb{A}))^{-\frac{p}{2}}\|\\
&\le \frac{1}{4}\|k^{\frac{p}{4}}\Phi(\sum\limits_{i=1}^nw_iA_i)^\frac{p}{2}+k^{-\frac{p}{4}}M^\frac{p}{2}m^\frac{p}{2}\Phi(P_t(\omega;\mathbb{A}))^{-\frac{p}{2}}\|^2~~(\text{by}~ \eqref{lem12})\\
&\le \frac{1}{4}\|k^{\frac{\alpha}{2}}\Phi(\sum\limits_{i=1}^nw_iA_i)^\alpha+k^{-\frac{\alpha}{2}}M^\alpha m^\alpha\Phi(P_t(\omega;\mathbb{A}))^{-\alpha}\|^\frac{p}{\alpha}~~(\mathrm{by}~ \eqref{lem2})\\
&\le \frac{1}{4}\|k^{\frac{\alpha}{2}}\Phi(\sum\limits_{i=1}^nw_iA_i)^\alpha+k^{\frac{\alpha}{2}}M^\alpha m^\alpha\Phi(\sum\limits_{i=1}^nw_iA_i)^{-\alpha}\|^\frac{p}{\alpha}~~(\mathrm{by}~ \rm \eqref{th343})\\
&\le \frac{k^{\frac{p}{2}}(M^\alpha+m^\alpha)^\frac{p}{\alpha}}{4}.
\end{align*}
The last inequality above holds as follows:  The condition $m\le A_i\le M$ implies that
$m^\alpha\le\Phi(\sum\limits_{i=1}^nw_iA_i)^\alpha\le M^\alpha$.
Therefore, the following inequalty
\[
\Phi(\sum\limits_{i=1}^nw_iA_i)^\alpha+M^\alpha m^\alpha\Phi(\sum\limits_{i=1}^nw_iA_i)^{-\alpha}\le M^\alpha+m^\alpha
\]
holds for any positive unital linear map $\Phi$ and a weight $\omega=(w_1,\cdots,w_n)\in \Delta_n$. Thus, \eqref{th341} holds.

By a similar argument, \eqref{th342} can be derived from the inequality \eqref{rem332}.
\end{proof}

Taking $\alpha=2$ in Theorem 3.4, we have
\begin{corollary}
Under the same conditions as in Theorem \ref{th34}, then for $p\ge 4$,
\begin{equation}
\label{cor341}
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(k(M^2+m^2))^{p}}{16M^pm^p} \Phi(P_t(\omega;\mathbb{A}))^p
\end{equation}
and
\begin{equation}
\label{cor342}
\Phi(\sum\limits_{i=1}^nw_iA_i)^p\le \frac{(k(M^2+m^2))^{p}}{16M^pm^p} P_t(\omega;\Phi(\mathbb{A}))^p.
\end{equation}
\end{corollary}
\begin{remark}
If $\frac{M}{m}\le 2+\sqrt{3}$, we have $k(M^2+m^2)\le (M+m)^2$, so \eqref{cor341} and \eqref{cor342} are sharper than \eqref{th341}
and \eqref{th342} for $p\ge 4$, respectively.
\end{remark}

Note that we can also generalize the inequality \eqref{th341}, \eqref{th342}, \eqref{cor341} and \eqref{cor342} to the Karcher mean by \eqref{Lim} similarly.
\section{Weighted arithmetic and geometric means due to Lawson and Lim}
In 2006, Yamazaki \cite{Yamazaki} obtained a converse of arithmetic-geometric means inequality of $n$-operators via Kantorovich constant by induction on $n$. Soon after, Fujii el al. \cite{Fujii} also proved a stronger reverse inequality of the weighted arithmetic and geometric means due to Lawson and Lim of $n$-operators by the Kantorovich inequality in Theorem \ref{th21}.

In this section, we obtain the higher-power reverse inequalities of the weighted arithmetic and geometric means due to Lawson and Lim of $n$-operators, and several complements of the weighted geometric mean for $n$-variables have been established.
\begin{theorem}\label{th41}
For any positive integer $n\ge 2$, let $A_1, A_2, \cdots, A_n$ be positive
 invertible operators on a Hilbert space $\mathcal{H}$ such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then for $p\ge 2$,
\begin{equation}\label{th411}
A[n,t](A_1,\cdots, A_n)^p\le \frac{(m+M)^{2p}}{16m^pM^p}G[n,t](A_1,\cdots, A_n)^p.
\end{equation}
\end{theorem}

\begin{proof}
Let a map $\Psi: \mathcal{B}(\mathcal{H})\oplus \cdots \oplus\mathcal{B}(\mathcal{H})
\mapsto \mathcal{B}(\mathcal{H})\oplus \cdots \oplus\mathcal{B}(\mathcal{H})$ be defined by
\[\Psi\left(
  \begin{array}{ccc}
    A_1  & &0 \\
     & \ddots& \\
     0& &A_n\\
  \end{array}
\right)=
\left(
  \begin{array}{ccc}
    t[n]_1A_1+\cdots+t[n]_nA_n  & &0 \\
     & \ddots& \\
     0& &t[n]_1A_1+\cdots+t[n]_nA_n\\
  \end{array}
\right).
\]
Then $\Psi$ is a positive linear map such that $\Psi(I)=I$.
The condition $0<m\le A_i\le M$ for $i=1,2,\cdots,n$ implies that
\begin{equation}\label{th412}
m\left(
  \begin{array}{ccc}
    I  & &0 \\
     & \ddots& \\
     0& &I\\
  \end{array}
\right)\le
\Psi\left(
  \begin{array}{ccc}
    A_1  & &0 \\
     & \ddots& \\
     0& &A_n\\
  \end{array}
\right)
\le M\left(
  \begin{array}{ccc}
    I  & &0 \\
     & \ddots& \\
     0& &I\\
  \end{array}
\right).
\end{equation}
By (2.3) in \cite{Lin2}, we have
\[
\Psi\left(
  \begin{array}{ccc}
    A_1  & &0 \\
     & \ddots& \\
     0& &A_n\\
  \end{array}
\right)+
Mm\Psi\left(
  \begin{array}{ccc}
    A_1^{-1}  & &0 \\
     & \ddots& \\
     0& &A_n^{-1}\\
  \end{array}
\right)
\le (M+m)\left(
  \begin{array}{ccc}
    I  & &0 \\
     & \ddots& \\
     0& &I\\
  \end{array}
\right).
\]
Thus,
\begin{equation}
\label{th413}
A[n,t](A_1,\cdots, A_n)+Mm A[n,t](A_1^{-1},\cdots, A_n^{-1})\le m+M.
\end{equation}
On the other hand, by computing, we deduce
\begin{align*}
\|M^\frac{p}{2}m^\frac{p}{2}&A[n,t](A_1,\cdots, A_n)^\frac{p}{2}G[n,t](A_1,\cdots, A_n)^{-\frac{p}{2}}\|\\
&\le \frac{1}{4}\|A[n,t](A_1,\cdots, A_n)^\frac{p}{2}+M^\frac{p}{2}m^\frac{p}{2}G[n,t](A_1,\cdots, A_n)^{-\frac{p}{2}}\|^2~(\text{by}~ \eqref{lem12})\\
&\le \frac{1}{4}\|A[n,t](A_1,\cdots, A_n)+MmG[n,t](A_1,\cdots, A_n)^{-1}\|^p~(\text{by}~ \eqref{lem2})\\
&= \frac{1}{4}\|A[n,t](A_1,\cdots, A_n)+MmG[n,t](A_1^{-1},\cdots, A_n^{-1})\|^p\\
&\le \frac{1}{4}\|A[n,t](A_1,\cdots, A_n)+MmA[n,t](A_1^{-1},\cdots, A_n^{-1})\|^p~(\text{by}~ \eqref{agh})\\
&\le \frac{(M+m)^p}{4}.~(\text{by}~ \eqref{th413})
\end{align*}

The equality above follows from the self-duality of the geometric mean (See \cite{Ando1, Fujii, Yamazaki}).
\end{proof}
Taking $p=2$,  \eqref{th411} implies the following corollary:
\begin{corollary}\label{cor41}
For any positive integer $n\ge 2$, let $A_1, A_2, \cdots, A_n$ be positive
 invertible operators on a Hilbert space $\mathcal{H}$ such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then
\begin{equation}\label{cor411}
A[n,t](A_1,\cdots, A_n)^2\le \frac{(m+M)^{4}}{16m^2M^2}G[n,t](A_1,\cdots, A_n)^2.
\end{equation}
\end{corollary}

Note that if $t=\frac{1}{2}$, the inequality \eqref{cor411} reduces to Lin's result (See \cite[Theorem 3.2]{Lin1}).
\begin{theorem}\label{th42}
For any positive integer $n\ge 2$, let $A_1, A_2, \cdots, A_n$ be positive
 invertible operators on a Hilbert space $\mathcal{H}$ such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then for $1<\alpha\le 2$ and $p\ge2\alpha$,
\begin{equation}\label{th421}
A[n,t](A_1,\cdots, A_n)^p\le \frac{(k^{\frac{\alpha}{2}}(M^\alpha+m^\alpha))^{\frac{2p}{\alpha}}}{16M^pm^p}G[n,t](A_1,\cdots, A_n)^p,
\end{equation}
where $k=\frac{(m+M)^{2}}{4mM}$.
\end{theorem}
\begin{proof}
Let a map $\Psi: \mathcal{B}(\mathcal{H})\oplus \cdots \oplus\mathcal{B}(\mathcal{H})
\mapsto \mathcal{B}(\mathcal{H})\oplus \cdots \oplus\mathcal{B}(\mathcal{H})$ be defined as in the proof of Theorem \ref{th41}.
By \eqref{th412}, we have
\[
m^\alpha\left(
  \begin{array}{ccc}
    I  & &0 \\
     & \ddots& \\
     0& &I\\
  \end{array}
\right)\le
\Psi^\alpha\left(
  \begin{array}{ccc}
    A_1  & &0 \\
     & \ddots& \\
     0& &A_n\\
  \end{array}
\right)
\le M^\alpha\left(
  \begin{array}{ccc}
    I  & &0 \\
     & \ddots& \\
     0& &I\\
  \end{array}
\right),
\]
that is,
\[m^\alpha\le A[n,t](A_1,\cdots, A_n)^\alpha\le M^\alpha.\]
By (2.3) in \cite{Lin2}, we have
\begin{equation}\label{th423}
A[n,t](A_1,\cdots, A_n)^\alpha+M^\alpha m^\alpha A[n,t](A_1,\cdots, A_n)^{-\alpha}\le m^\alpha+M^\alpha.
\end{equation}
On the other hand, by \eqref{cor411},
\begin{equation}\label{th424}
k^{-\alpha}G[n,t](A_1,\cdots, A_n)^{-\alpha}\le A[n,t](A_1,\cdots, A_n)^{-\alpha}.
\end{equation}
By computing, we deduce
\begin{align*}
\|k^{-\frac{p}{2}}m^{\frac{p}{2}}M^{\frac{p}{2}}&A[n,t](A_1,\cdots, A_n)^\frac{p}{2}G[n,t](A_1,\cdots, A_n)^{-\frac{p}{2}}\|\\
&\le \frac{1}{4}\|A[n,t](A_1,\cdots, A_n)^\frac{p}{2}+k^{-\frac{p}{2}}m^{\frac{p}{2}}M^{\frac{p}{2}}G[n,t](A_1,\cdots, A_n)^{-\frac{p}{2}}\|^2~(\text{by}~ \eqref{lem12})\\
&\le \frac{1}{4}\|(A[n,t](A_1,\cdots, A_n)^\alpha+k^{-\alpha}m^{\alpha}M^{\alpha}G[n,t](A_1,\cdots, A_n)^{-\alpha})^{\frac{p}{2\alpha}}\|^2~(\text{by}~ \eqref{lem2})\\
&= \frac{1}{4}\|A[n,t](A_1,\cdots, A_n)^\alpha+k^{-\alpha}m^{\alpha}M^{\alpha}G[n,t](A_1,\cdots, A_n)^{-\alpha}\|^{\frac{p}{\alpha}}\\
&\le \frac{1}{4}\|A[n,t](A_1,\cdots, A_n)^\alpha+m^{\alpha}M^{\alpha}A[n,t](A_1,\cdots, A_n)^{-\alpha}\|^p~(\text{by}~ \eqref{th424})\\
&\le \frac{(M^{\alpha}+m^{\alpha})^p}{4}.~(\text{by}~ \eqref{th423})
\end{align*}
We obtain the desired result.
\end{proof}

Putting $\alpha=2$ in the inequality \eqref{th421}, which implies that
\begin{corollary}
Under the same conditions as in Theorem \ref{th42}, then for $p\ge 4$,
\begin{equation}
\label{cor421}
A[n,t](A_1,\cdots, A_n)^p\le \frac{(k(M^2+m^2))^{p}}{16M^pm^p}G[n,t](A_1,\cdots, A_n)^p.
\end{equation}
\end{corollary}
\begin{remark}

When $\frac{M}{m}\le 2+\sqrt{3}$, it is easy to see that \eqref{cor421} is sharper than \eqref{th411} for $p\ge 4$.
\end{remark}

Next, we show the complements of the weighted geometric mean due to Lawson and Lim by virtue of the following lemma.
\begin{lemma}
For any integer $n\ge 2$, let $A_1,A_2,\cdots,A_n$ be positive
 invertible operators in $\mathbb{P}$
such that $m\le A_i\le M$ for all $i=1,2,\cdots,n$ and some scalars $0<m\le M$.
Then
\begin{equation}
\label{lem411}
A[n,t](A^p_1,\cdots,A^p_n)\le K(m,M,p)A[n,t](A_1,\cdots, A_n)^p~for~all~p\ge1,
\end{equation}
where $K(m,M,p)=\frac{(p-1)^{p-1}}{p^p}\frac{(M^p-m^p)^p}{(M-m)(mM^p-Mm^p)^{p-1}}$ is the generalized Kantorovich constant.
\end{lemma}
\begin{proof}
By Corollary 2.6 in \cite{Micic},
\[\Phi(A^p)\le K(m,M,p)\Phi(A)^p~for~all~p\ge1.\]
Let the map $\Phi: \mathcal{B}(\mathcal{H})\oplus \cdots \oplus\mathcal{B}(\mathcal{H})
\mapsto \mathcal{B}(\mathcal{H})\oplus \cdots \oplus\mathcal{B}(\mathcal{H})$ be defined as $\Psi$ in the proof of Theorem \ref{th41}. Then for $p\ge1$,
 \[A[n,t](A^p_1,\cdots,A^p_n)\le K(m,M,p)A[n,t](A_1,\cdots, A_n)^p.\qedhere\]
\end{proof}

\begin{theorem} \label{th43}
For any integer $n\ge 2$, let $A_1,A_2,\cdots,A_n$ be positive invertible operators in $\mathbb{P}$
such that $m\le A_i\le M $ for all $i=1,2,\cdots,n$ and some scalars $0<m\le M$.
Then
\[
G[n,t](A^p_1,\cdots,A^p_n)\le K(m,M,p)\frac{(m+M)^{2p}}{4^pm^pM^p}G[n,t](A_1,\cdots, A_n)^p~~for~all~1<p\le2,
\]
and
\[
~G[n,t](A^p_1,\cdots,A^p_n)\le K(m,M,p)\frac{(m+M)^{2p}}{16m^pM^p}G[n,t](A_1,\cdots, A_n)^p~for~all~p\ge2.~~~~~~
\]
\end{theorem}
\begin{proof}
By the arithmetic-geometric mean inequality and \eqref{lem411}, it follows that
\begin{equation}
\label{th431}
G[n,t](A^p_1,\cdots,A^p_n)\le A[n,t](A^p_1,\cdots,A^p_n)\le K(m,M,p)A[n,t](A_1,\cdots, A_n)^p~for~p\ge1.
\end{equation}

For $p\in(1,2]$, it follows from \eqref{cor411} and the L\"{o}ewner-Heinz inequality that
\[A[n,t](A_1,\cdots, A_n)^p\le \left(\frac{(m+M)^{2}}{4mM}\right)^pG[n,t](A_1,\cdots, A_n)^p.\]
Combining these inequalities above, we have
\[G[n,t](A^p_1,\cdots,A^p_n)\le K(m,M,p)\frac{(m+M)^{2p}}{4^pm^pM^p}G[n,t](A_1,\cdots, A_n)^p.\]

For $p\in[2,\infty)$, from \eqref{th411} and \eqref{th431}, we obtain
\[G[n,t](A^p_1,\cdots,A^p_n)\le K(m,M,p)\frac{(m+M)^{2p}}{16m^pM^p}G[n,t](A_1,\cdots, A_n)^p.\qedhere\]
\end{proof}

In the following remark, we present the case of $p\ge 1$ about \eqref{th232} for the Ando-Li-Mahthias geometric mean.
\begin{remark}
Let $t=\frac{1}{2}$ in Theorem \ref{th43}. Then
\[
G_{ALM}(A^p_1,\cdots,A^p_n)\le K(m,M,p)\frac{(m+M)^{2p}}{4^pm^pM^p}G_{ALM}(A_1,\cdots, A_n)^p~~for~all~1<p\le2
\]
and
\[
~G_{ALM}(A^p_1,\cdots,A^p_n)\le K(m,M,p)\frac{(m+M)^{2p}}{16m^pM^p}G_{ALM}(A_1,\cdots, A_n)^p~~for~all~p\ge2.~~~~~
\]
\end{remark}

\begin{theorem} \label{th44}
For any integer $n\ge 2$, let $A_1,A_2,\cdots,A_n$ be positive invertible operators  in $\mathbb{P}$
such that $m\le A_i\le M $ for all $i=1,2,\cdots,n$ and some scalars $0<m\le M$.
Then
\[
G[n,t](A^p_1,\cdots,A^p_n)^\frac{1}{p}\le K\left(m^q,M^q,\frac{p}{q}\right)^\frac{1}{p}\left(\frac{(m^q+M^q)^{2}}{4m^qM^q}\right)^\frac{1}{q}G[n,t](A^q_1,\cdots,A^q_n)^\frac{1}{q}~~~
\]
for all $1<\frac{p}{q}\le2$ and $p\ge1$,
and
\[
~~G[n,t](A^p_1,\cdots,A^p_n)^\frac{1}{p}\le 4^{-\frac{2}{p}} K\left(m^q,M^q,\frac{p}{q}\right)^\frac{1}{p}\left(\frac{(m^q+M^q)^{2}}{m^qM^q}\right)^\frac{1}{q}G[n,t](A^q_1,\cdots,A^q_n)^\frac{1}{q}
\]
for all $\frac{p}{q}\ge2$ and $p\ge1$.
\end{theorem}
\begin{proof}
For each $0<q\le p$, it follows from the arithmetic-geometric mean inequality \eqref{agh} and \eqref{lem411} that
\begin{equation}
\label{th441}
\begin{split}
G[n,t](A^p_1,\cdots,A^p_n)&\le A[n,t](A^p_1,\cdots,A^p_n)\\
&= A[n,t]((A^q_1)^\frac{p}{q},\cdots,(A^q_n)^\frac{p}{q})\\
&\le K\left(m^q,M^q,\frac{p}{q}\right)A[n,t](A^q_1,\cdots, A^q_n)^\frac{p}{q}.
\end{split}
\end{equation}
On the other hand, for $1<\frac{p}{q}\le2$, from \eqref{cor411} and $m^qI\le A_i^q\le M^qI$, it follows that
\[
A[n,t](A^q_1,\cdots, A^q_n)^\frac{p}{q}\le \left(\frac{(m^q+M^q)^{2}}{4m^qM^q}\right)^\frac{p}{q}G[n,t](A^q_1,\cdots,A^q_n)^\frac{p}{q}.
\]
Combining the two inequalities above, we obtain
\[
G[n,t](A^p_1,\cdots,A^p_n)\le K\left(m^q,M^q,\frac{p}{q}\right)\left(\frac{(m^q+M^q)^{2}}{4m^qM^q}\right)^\frac{p}{q}G[n,t](A^q_1,\cdots,A^q_n)^\frac{p}{q}.
\]
By the L\"{o}ewner-Heinz inequality and $p\ge1$, it follows that
\[
G[n,t](A^p_1,\cdots,A^p_n)^\frac{1}{p}\le K\left(m^q,M^q,\frac{p}{q}\right)^\frac{1}{p}\left(\frac{(m^q+M^q)^{2}}{4m^qM^q}\right)^\frac{1}{q}G[n,t](A^q_1,\cdots,A^q_n)^\frac{1}{q}.
\]
Similarly, for all $\frac{p}{q}\ge2$, from \eqref{th411} we have
\[
A[n,t](A^q_1,\cdots, A^q_n)^\frac{p}{q}\le 4^{-2}\left(\frac{(m^q+M^q)^{2}}{m^qM^q}\right)^\frac{p}{q}G[n,t](A^q_1,\cdots,A^q_n)^\frac{p}{q}.
\]
Combining with \eqref{th441}, we obtain
\[
G[n,t](A^p_1,\cdots,A^p_n)\le 4^{-2}K\left(m^q,M^q,\frac{p}{q}\right)\left(\frac{(m^q+M^q)^{2}}{m^qM^q}\right)^\frac{p}{q}G[n,t](A^q_1,\cdots,A^q_n)^\frac{p}{q}.
\]
It follows from $p\ge1$ that
\[
G[n,t](A^p_1,\cdots,A^p_n)^\frac{1}{p}\le 4^{-\frac{2}{p}}K\left(m^q,M^q,\frac{p}{q}\right)^\frac{1}{p}\left(\frac{(m^q+M^q)^{2}}{m^qM^q}\right)^\frac{1}{q}G[n,t](A^q_1,\cdots,A^q_n)^\frac{1}{q}.
\]
This completes the proof.
\end{proof}
\begin{remark}
Although $\lim_{q\rightarrow 0}G[n,t](A^q_1,\cdots,A^q_n)^\frac{1}{q}=\diamondsuit(\hat{\omega}; A_1,\cdots, A_n)$ (the chaotic geometric mean),
$\lim_{q\rightarrow 0}\left(\frac{(m^q+M^q)^{2}}{4m^qM^q}\right)^\frac{1}{q}=1$ and $\lim_{q\rightarrow 0}K\left(m^q,M^q,\frac{p}{q}\right)^\frac{1}{p}=S(h^p)^\frac{1}{p}$ (See \cite{Fujii1}), we can not obtain
\[
G[n,t](A^p_1,\cdots,A^p_n)^\frac{1}{p}\le S(h^p)^\frac{1}{p}\diamondsuit(\hat{\omega}; A_1,\cdots, A_n)
\]
for all $p\ge1$ from Theorem \ref{th44}.
\end{remark}

\section{Comparisons between the weighted Karcher mean and the Lawson-Lim geometric mean}
In the final section, we make comparisons between the weighted Karcher mean and the Lawson-Lim geometric mean for higher power.
This is a fascinating work because the order relation can be preserved between higher power operators by the Kantorovich constant.
\begin{lemma}\label{lem51}\cite{Fujii2}
Let $0<m\le A\le M$ and $A\le B$. Then
\[
A^2\le \frac{(M+m)^2}{4Mm}B^2.
\]
\end{lemma}

\begin{lemma}\label{lem52}\cite{Furuta1}
Let $A$ and $B$ be positive invertible operators on a Hilbert space $\mathcal{H}$ satisfying
$B\ge A>0$ and $0<m\le A\le M$. Then
\[
\left(\frac{M}{m}\right)^{p-1}B^p\ge\mathrm{K}(m,M,p)B^p\ge A^p
\]
holds for any $p\ge 1$, where $\mathrm{K}(m,M,p)$ is the generalized Kantorovich constant or the Ky Fan-Furuta constant and
\[\left(\frac{M}{m}\right)^{p-1}\ge \mathrm{K}(m,M,p).\]
\end{lemma}
\begin{theorem}
For any positive integer $n\ge 2$, let $A_1, A_2, \cdots, A_n$ be positive
invertible operators on a Hilbert space $\mathcal{H}$ such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then
\begin{equation}
\label{th511}
\begin{split}
\left(\frac{(m+M)^{2}}{4mM}\right)^{-3}G[n,t](A_1,\cdots, A_n)^2&\le G_K(\hat{\omega}, A_1,\cdots, A_n)^2\\
&\le \left(\frac{(m+M)^{2}}{4mM}\right)^3G[n,t](A_1,\cdots, A_n)^2,
\end{split}
\end{equation}
where $\hat{\omega}=\left(t[n]_1, t[n]_2, \cdots, t[n]_n\right)$. 
\end{theorem}
\begin{proof}
The first inequality in \eqref{th511} follows from Lemma \ref{lem51}, the arithmetic-geometric mean inequality \eqref{agh} and \eqref{rem351} with $p=2$ and an identity map $\Phi$ that
\begin{align*}
G[n,t](A_1,\cdots, A_n)^2&\le \frac{(m+M)^{2}}{4mM}A[n,t](A_1,\cdots, A_n)^2\\
&\le\frac{(m+M)^{2}}{4mM}\left(\frac{(m+M)^{2}}{4mM}\right)^2 G_K(\hat{\omega}, A_1,\cdots, A_n)^2.
\end{align*}
The second inequality in \eqref{th511} follows from Lemma \ref{lem51} and \eqref{cor411} that
\begin{align*}
G_K(\hat{\omega}, A_1,\cdots, A_n)^2&\le \frac{(m+M)^{2}}{4mM}A[n,t](A_1,\cdots, A_n)^2\\
&\le\frac{(m+M)^{2}}{4mM}\left(\frac{(m+M)^{2}}{4mM}\right)^2 G[n,t]A_1,\cdots, A_n)^2.\qedhere
\end{align*}
\end{proof}

\begin{theorem}
For any positive integer $n\ge 2$, let $A_1, A_2, \cdots, A_n$ be positive
invertible operators on a Hilbert space $\mathcal{H}$ such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then
\begin{equation}
\label{th521}
G_K(\hat{\omega},A_1,\cdots, A_n)^p\le \mathrm{K}(m,M,p)\frac{(M+m)^{2p}}{4^pM^pm^p}G[n,t](A_1,\cdots, A_n)^p ~for ~all~ 1\le p\le 2
\end{equation}
and
\begin{equation}
\label{th522}
G_K(\hat{\omega},A_1,\cdots, A_n)^p\le \mathrm{K}(m,M,p)\frac{(M+m)^{2p}}{16M^pm^p}G[n,t](A_1,\cdots, A_n)^p ~for~ all~ p\ge 2.~~~~
\end{equation}
\end{theorem}

\begin{proof}
By the L\"{o}ewner-Heinz inequality and \eqref{cor411}, we have
\begin{equation}
\label{th523}
A[n,t](A_1,\cdots, A_n)^p\le\frac{(M+m)^{2p}}{4^pM^pm^p}G[n,t](A_1,\cdots, A_n)^p ~for ~all~ 1\le p\le 2.
\end{equation}
It follows from Lemma \ref{lem52} and \eqref{th523} that
\begin{align*}
G_K(\hat{\omega},A_1,\cdots, A_n)^p&\le \mathrm{K}(m,M,p)A[n,t](A_1,\cdots, A_n)^p\\
&\le\mathrm{K}(m,M,p)\frac{(M+m)^{2p}}{4^pM^pm^p}G[n,t](A_1,\cdots, A_n)^p.
\end{align*}
The inequality \eqref{th522} follows from Lemma \ref{lem52} and \eqref{th411} that
\begin{align*}
G_K(\hat{\omega},A_1,\cdots, A_n)^p&\le \mathrm{K}(m,M,p)A[n,t](A_1,\cdots, A_n)^p\\
&\le\mathrm{K}(m,M,p)\frac{(M+m)^{2p}}{16M^pm^p}G[n,t](A_1,\cdots, A_n)^p.\qedhere
\end{align*}
\end{proof}

\begin{remark}
When $p=2$, the inequality \eqref{th521} is equivalent to the second one in \eqref{th511}. So
\eqref{th521} generalize the second inequality in \eqref{th511}.
When $p=1, t=\frac{1}{2}$, the inequality \eqref{th521} reduces to the first one in \eqref{th231}.
\end{remark}
Next, we make a comparison of the weighted Karcher mean and the Lawson-Lim geometric mean for unitarily invariant norm by virtue of the following lemma.
\begin{lemma}\label{lem53}\cite[Lemma 3.2]{Fujii1}
Let $A$ and $B$ be positive invertible operators on a Hilbert space $\mathcal{H}$. If $A\le B$,
then there exists a unitary operator $U$ such that $A^p\le UB^pU^*$ for all $p>0$.
\end{lemma}

\begin{theorem}
For any positive integer $n\ge 2$, let $A_1, A_2, \cdots, A_n$ be positive invertible operators on a Hilbert space $\mathcal{H}$  such that $m\le A_i\le M$ for $i=1,2,\cdots,n$
and some scalars $0<m<M$. Then
\[|||G_K(\hat{\omega}, A_1,\cdots, A_n)^p|||\le \mathrm{K}(m,M,p)\frac{(M+m)^{2p}}{4^pM^pm^p}|||G[n,t](A_1,\cdots, A_n)^p|||\]
for $p\ge 1$ and every unitarily invariant norm $|||\cdot|||$.
\end{theorem}

\begin{proof}
By Theorem \ref{th21} and Lemma \ref{lem53}, there exists a unitary operator $U$ such that
\[A[n,t](A_1,\cdots, A_n)^p\le\left(\frac{(m+M)^{2}}{4mM}\right)^pUG[n,t](A_1,\cdots, A_n)^pU^*.\]
Combining Lemma \ref{lem52} with the above inequality, we have
\begin{align*}
G_K(\hat{\omega},A_1,\cdots, A_n)^p&\le \mathrm{K}(m,M,p)A[n,t](A_1,\cdots, A_n)^p\\
&\le\mathrm{K}(m,M,p)\left(\frac{(m+M)^{2}}{4mM}\right)^pUG[n,t](A_1,\cdots, A_n)^pU^*.\qedhere
\end{align*}
\end{proof}


\bibliographystyle{amsplain}

\end{document}